OpenStudy (anonymous):

Need help check a answer What is the perimeter of a triangle with the given side lengths? s1 = x + 4 cm s2 = 3x + 1 cm s3 = 7x + 3 cm Is it (11x)+(7)?

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

ok, perimeter you add them right?

OpenStudy (anonymous):

That's what did, I hope I'm right

OpenStudy (mathstudent55):

The x part is correct. Check the numbers part.

OpenStudy (anonymous):

yes, now add the x's first. x+3x+7x=? Now the numbers:4+1+3=?

OpenStudy (mathstudent55):

x + 3x + 7x = 11x since x is the same as 1x. What is 4 + 1 + 3?

OpenStudy (anonymous):

4 and x aren't the same

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

well then why 1x + 4 + 3?

OpenStudy (anonymous):

Oh wait nvm

OpenStudy (anonymous):

I forgot the other 1 sorry

OpenStudy (anonymous):

so is it (11x)+(8)?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

kk I'll wait till mathstudent is done

OpenStudy (mathstudent55):

We are adding these three lengths: \((x + 4) +(3x + 1) + (7x + 3) \) Since we only have additions, we can drop the parentheses to give us: \(=x + 4 +3x + 1 + 7x + 3 \) Using the commutative and associative properties, we can change the order and grouping: \(= x+ 3x + 7x + 4 + 1 + 3 \) \(= (x+ 3x + 7x) + (4 + 1 + 3) \) Now we combine like terms. Remember that x is the same as 1x, so we get: \(= 11x + 8\)

OpenStudy (anonymous):

Ah, ok thanks mathstudent! I'll fan you

OpenStudy (mathstudent55):

@LoveIt You are correct. You don't need the parentheses, but your answer is still correct.

OpenStudy (anonymous):

kk, ty!

OpenStudy (anonymous):

cya

OpenStudy (anonymous):

YAY I GOT 100%

OpenStudy (anonymous):

Thanks to you guys!

OpenStudy (anonymous):

your welcome!